# Anything Goes? A Crosslinguistic Study of (Im)possible Language Learning in LMs

### v1

@@ 4.2 Results & Discussion
Results
Our results are visualized in Figure 3 (bottom subgraph). Surprisingly, all five NP-perturbed languages exhibit lower perplexity compared to their attested counterparts across all four languages. Two of these (Nnda and anNd) are unattested in typological studies and are ruled out by generative approaches (Cinque, 2005), but we do not observe a significant difference in perplexity between the three languages with attested NP orders and the two languages with unattested orders.
When these POS tags are shuffled within NPs, perplexity increases, reaching or exceeding the perplexity of the attested languages. This rules out the possibility that limited perturbations simply do not affect model training.
To summarize, this experiment shows that, unlike humans, language models fail to show a gradient of difficulty in learning different NP orders according to their typological prevalence.
Discussion: Why Can’t LMs Distinguish Between Attested and Unattested Languages?
@@ Results
To make sense of these results, we propose two key factors that influence LM learning outcomes: randomness and constituency structure. By randomness, we refer to whether the perturbation function produces a perturbed text that can be deterministically recovered to its original form. By constituency structure, we mean whether the phrase structures of the original language are preserved in the perturbed version.
@@
Regarding randomness, string distributions with higher entropy are always harder to learn. This explains why NP-perturbed unattested languages show lower perplexity than attested languages and np_random variants. The reasoning is that our perturbation procedure enforces a strict ordering procedure, which may be (sometimes) violated in the original attested language. For example, although English is a dnaN language, certain constructions such as the dann (det-adj-num-noun; e.g., a beautiful five days) does not follow the dominant pattern. Once POS tag orders are normalized within NPs, the resulting constructions become more predictable. Therefore, all normalized NPs, including our unattested NPs, may have lower overall entropy, which could explain why they are easier to learn. In fact, the normalized dnaN, which has the same typical word order as English, shows lower perplexity than the original, unnormalized English; and the same applies to our other languages in this experiment.
Regarding constituency structure, we hypothesize that disrupting constituency weakens local dependency relations within phrase structures. This explains why all impossible languages in the previous experiments, despite maintaining a deterministic order, still results in higher perplexity than NP-perturbed languages (Figure 3). Similarly, this may also explain the higher perplexity of count-based grammars in Kallini et al. (2024): the insertion of morphological markers disrupts phrase structure integrity. One exception is reverse_full, which preserves constituency structure while maintaining a deterministic order. As shown in Figure 3, reverse_full exhibits perplexity closer to unattested languages. We hypothesize that this exception may be due to other factors, such as information density (Clark et al., 2023). Since reversing word order alters information flow, it may obscure more accessible information. For example, predicting a pronoun given a preceding noun is easier than predicting a noun given a preceding pronoun, potentially increasing difficulty for the LM. One shortcoming of this experiment is that it was based entirely on NP perturbations(Kallini et al., 2024). Future work could investigate whether similar effects occur when POS order or constituency structure is disrupted on a larger scale.
7 Conclusion
In this paper, we extend Kallini et al. (2024) to a broader multilingual context. Our experiments provide mixed results that complement those of Kallini et al. (2024). We find that while GPT-2 small can distinguish between attested languages and their impossible variants, its learning outcomes do not separate all attested and all unattested or impossible languages. That being said, LMs do tend to learn attested languages better, on average, than impossible languages, and we achieve a separability of $0.75$ between the two classes based on the models’ perplexity. Finally, we observe that some unattested languages show lower perplexity than their attested counterparts even though they exhibit NP orderings that flout Greenberg’s Universal 20.
@@ 7 Conclusion
What to make of these results in the context of our original question–whether LLMs can serve as cognitive models? While our results show that GPT-2 does not behave as we might expect from a fully human-like learner, they also demonstrate that it has a soft preference for attested over impossible languages. Skeptics have previously likened LLMs to a bad theory of physics in which “anything goes.” [^10] Chomsky, quoted from an email to Gary Marcus: You can’t go to a physics conference and say: I’ve got a great theory. It accounts for everything and is so simple it can be captured in two words: “Anything goes.” All known and unknown laws of nature are accommodated, no failures. Of course, everything impossible is accommodated also. In line with Kallini et al. (2024), our results demonstrate that these models do not instantiate an “anything goes” hypothesis. Rather, their incremental data-processing architectures represent a useful starting point for studying human language processing. Refining models to achieve stronger alignment with people is possible, and will likely lead to lasting insights about human cognitive architecture.
@@ 8 Limitations
We acknowledge that our experiments rely on GPT-2 Small, which may not generalize to larger models. This choice was made for two reasons: (1) running experiments across multiple languages is computationally expensive; (2) we aimed for comparability with Kallini et al. (2024). Future work could explore whether our findings hold for larger models. Additionally, the dataset used for training the language model is relatively small. This is a deliberate trade-off between data size and linguistic diversity. While a larger dataset might yield more robust results, our approach ensures broader typological coverage. Lastly, in our experiments on unattested languages, we generated synthetic data by perturbing languages based on Universal 20. However, linguistic correlations extend beyond word order universals. For instance, Greenbergian correlations (Dryer, 1992) suggest that verb-object order often correlates with other features such as adposition-noun phrase order and determiner-noun phrase order. Future work will explore more nuanced perturbations to better capture such cross-linguistic dependencies.

### v2

@@ 4.2 Results & Discussion
Figure 4: Mean $\Delta$ GenScore across four languages under five NP perturbations. Error bars indicate the 95% CI computed over three random seeds. Positive $\Delta$ GenScore indicates better generalization for models trained on attested languages, while negative values indicate better generalization for models trained on unattested languages.
Studies such as Xu et al. (2025) suggest a difference between models’ perplexity and results of targeted evaluations. Therefore, we additionally conduct a targeted test to assess how well LMs trained on different perturbed languages generalize. Specifically, we propose $\Delta$GenScore to quantify their generalization ability, measured across a test corpus of $n$ sentences, and defined as:
[…]
If a model assigns higher probability to natural (i.e., attested) word orders regardless of its training data, then it would obtain a $\Delta$GenScore of $1$. Likewise, if it assigns a higher probability to unattested orders regardless of its training data, then it would have a $\Delta$GenScore of $-1$. A $\Delta$GenScore of 00 indicates that the model always assigns higher probabilities to sequences that match its training data. Therefore, we interpret positive $\Delta$GenScore values as indicating better generalization for natural word orderings, and negative scores as indicating better generalization for perturbed orderings. We use $\Delta$GenScore to investigate models trained on each of our natural languages, and compare them to each of our possible NP perturbations.
6.1 Results
Perplexity
Our results (Figure 3, bottom subgraph) show that shuffling POS tags within NPs increases perplexity, often matching or exceeding that of attested languages. This rules out the possibility that limited perturbations do not affect model training. Surprisingly, all five NP-perturbed languages exhibit lower perplexity than their attested counterparts across all four languages, though the differences are not significant for Italian, Chinese, and Portuguese (by a Welch’s t-test with Bonferroni correction). [^10] For English, there is no significant difference between shuffle_control and dnaN, the dominant NP order in English. No significant difference is observed between languages with attested (i.e., daNn, dnaN, and dnNa) and unattested NP orders (i.e., Nnda and anNd) either, indicating a lack of human alignment in language learning bias.
Generalization Test
The results from this experiment are visualized in Figure 4 and present a mixed picture. Two observations emerge. First, models trained on Nnda and anNd, the two typologically absent orderings, consistently yield positive $\Delta$GenScore across all languages. This indicates poorer generalization of models trained on unattested patterns than models on attested ones listed in Table 3. Second, $\Delta$GenScore remains positive for all five NP perturbations in English and Chinese but shows mixed results for Italian and Portuguese. Since English and Chinese predominantly follow the dnaN order and Italian and Portuguese follow dnNa, this suggests models trained on dnaN orders generalize more consistently. This finding, if confirmed, supports Culbertson and Newport (2015)’s report of human biases toward harmonic languages. However, for stronger conclusions, further investigation with more typologically diverse languages and NP perturbations is needed.
Summary
Experiment 3 shows that while LMs do not reflect a gradient of difficulty measured by perplexity in learning different NP orders based on typological prevalence, they may exhibit human-aligned generalization patterns for typologically unattested languages in the generalization test. The differences between perplexity and targeted evaluation results are consistent with Xu et al. (2025)’s findings, which show similar discrepancies.
6.2 Discussion of Perplexity Results
@@
Why doesn’t LM perplexity distinguish between attested and unattested languages? We propose two key factors that influence LM learning outcomes: randomness and constituency structure. By randomness, we refer to whether the perturbation function produces a perturbed text that can be deterministically recovered to its original form. By constituency structure, we mean whether the phrase structures of the original language are preserved in the perturbed version.
@@
Regarding randomness, as LMs are simply distributions over strings (Borenstein et al., 2024), introducing randomness increases unpredictability of the text, thus increasing the entropy of the sequence. This explains why NP-perturbed unattested languages show lower perplexity than attested languages and np_random variants. The reasoning is that our perturbation procedure enforces a strict ordering procedure, which may be (sometimes) violated in the original attested language. For example, although English is a dnaN language, certain constructions such as the dann (det-adj-num-noun; e.g., a beautiful five days) does not follow the dominant pattern. Once POS tag orders are normalized within NPs, the resulting constructions become more predictable. Therefore, all normalized NPs, including our unattested NPs, may have lower overall entropy, which could explain why they are easier to learn. In fact, the normalized dnaN, which has the same typical word order as English, shows lower perplexity than the original, unnormalized English; and the same applies to our other languages in this experiment.
Regarding constituency structure, we hypothesize that disrupting constituency weakens local dependency relations within phrase structures. This explains why in experiments 1 and 2, all LMs’ perplexities for impossible languages are higher than for NP-perturbed languages, despite maintaining a deterministic order (Figure 3). Similarly, this may also explain the higher perplexity of count-based grammars in Kallini et al. (2024), where the authors insert a morphological marker a certain number of words or tokens after a host word. The count-based insertion may disrupt phrase structure integrity. [^11] We do not replicate these count-based experiments.
In sum, this discussion points to a potential confound in our experiments: although the texts are parallel in content, languages with normalized NP structures may have lower entropy. In this case, even if LMs learn all languages equally well, lower entropy would naturally lead to lower perplexity. Future work could control for entropy across NP-perturbed languages to test whether perplexity differences persist.
7 Discussion & Conclusion
In this paper, we extend Kallini et al. (2024) to a broader multilingual context using two new parallel corpora. Our findings complement their work, suggesting that models have a preference for human-like languages, although the preference is somewhat gradient and depends on the testing setup. First, while GPT-2 small assigns lower perplexity to attested languages compared to their impossible variants, the difference is sometimes not significant, especially later in training. Second, the model does not fully separate all attested from all unattested or impossible languages, but it does generally learn attested languages better, achieving a separability of $0.75$ between the two classes based on perplexity. In the NP word order experiments, some unattested languages exhibit lower perplexity than their attested counterparts, despite having orderings that violate Greenberg’s Universal 20. However, when assessed using targeted evaluation methods, a more promising pattern emerges: GPT-2 seems to favor typologically attested, as opposed to unattested NP variants, and shows some preference for harmonic word orderings.
@@ 7 Discussion & Conclusion
What to make of these results in the context of our original question–whether LMs can serve as cognitive models? While our results show that GPT-2 does not behave as we might expect from a fully human-like learner, they also demonstrate that it has a soft preference for attested over impossible languages. Skeptics have previously linked LMs to a bad theory of physics in which “anything goes.” [^12] Chomsky, quoted from an email to Gary Marcus: “You can’t go to a physics conference and say: I’ve got a great theory. It accounts for everything and is so simple it can be captured in two words: ‘Anything goes.’ All known and unknown laws of nature are accommodated, no failures. Of course, everything impossible is accommodated also.” In line with Kallini et al. (2024), our results demonstrate that these models do not instantiate an “anything goes” hypothesis. Rather, their incremental data-processing architectures represent a useful starting point for studying human language processing and learning. Refining models to better align with humans is possible and will likely lead to lasting insights about human cognitive architecture.
@@ 8 Limitations
We acknowledge that our experiments rely on GPT-2 Small, which may not generalize to larger models. This choice was made for two reasons: (1) running experiments across multiple languages is computationally expensive; (2) we aimed for comparability with Kallini et al. (2024). Future work could explore whether our findings hold for larger models or similarly sized models with different architectures. Additionally, the dataset used for training the language model is relatively small. This is a deliberate trade-off between data size and linguistic diversity. While a larger dataset might yield more robust results, our approach ensures broader typological coverage. In our experiments on unattested languages, we generated synthetic data by perturbing languages based on Universal 20. However, linguistic correlations extend beyond word order universals. For instance, Greenbergian correlations (Dryer, 1992) suggest that verb-object order often correlates with other features such as adposition-noun phrase order and determiner-noun phrase order. Future work will explore more nuanced perturbations to better capture such cross-linguistic dependencies. Lastly, the data we use has not been manually checked yet. It is possible that our parallel corpora include noise that might influence the learning results.
@@ 9 Ethics Statement
10 Acknowledgments
We thank Amir Zeldes, Nathan Schneider, Yilun Zhu, Dan DeGenaro, Wesley Scivetti, and all members of PICoL and NERT for their helpful feedback and support.
@@ 10 Acknowledgments
In experiment 1, we also conducted a Spearman’s Ranking Correlation test between the results on OPUS30 English and those from Kallini et al. (2024)’s experiments. We grouped the shuffle_deterministic languages together and observed that the ranking of our English impossible variants aligns perfectly with that reported by Kallini et al. (2024) ($\rho=1$, $p=0.0027$).
@@ Appendix F POS tags of each treebank
Appendix G Statistical test between impossible languages
We conducted Welch’s paired t-test comparing different perturbations with shuffle_control across 12 checkpoints. The results are ordered alphabetically.
We find that for Dutch, Russian, and Turkish, the difference between shuffle_control and other perturbations is always significant; by contrast, for languages including Arabic, Chinese, English, German, and Romanian, the difference becomes less significant or insignificant in the locally shuffled variants.
Perturbation | Step 100 200 300 400 500 600 700 800 900 1000 1100 1200
perturb_reverse_full_word <0.001 <0.001 0.0036 0.0484 <0.001 0.0283 0.0781 0.5293 1 0.8811 1 1
shuffle_deterministic21 1 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_deterministic57 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_deterministic84 0.7871 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_even_odd <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_local10 0.4242 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_local2 1 <0.001 <0.001 <0.001 <0.001 0.7469 1 1 1 1 1 1
shuffle_local3 1 <0.001 <0.001 <0.001 <0.001 0.0075 0.0062 0.1182 1 1 1 1
shuffle_local5 0.4024 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 0.3782 0.3208 1 1
Table 8: Welch’s t-test comparing each perturbation with shuffle_control across 12 checkpoints for Arabic, with Bonferroni adjustment.
Perturbation | Step 100 200 300 400 500 600 700 800 900 1000 1100 1200
perturb_reverse_full_word <0.001 <0.001 <0.001 1 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_deterministic21 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_deterministic57 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_deterministic84 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_even_odd <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_local10 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_local2 <0.001 <0.001 <0.001 <0.001 <0.001 0.0142 1 1 1 1 1 1
shuffle_local3 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 0.0012 0.0014 0.0063 0.135
shuffle_local5 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 1 1 0.6033 0.0391 0.0016
Table 9: Welch’s t-test comparing each perturbation with shuffle_control across 12 checkpoints for Chinese, with Bonferroni adjustment.
Perturbation | Step 100 200 300 400 500 600 700 800 900 1000 1100 1200
perturb_reverse_full_word <0.001 <0.001 0.0012 0.1697 0.0396 0.0137 0.003 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_deterministic21 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_deterministic57 1 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_deterministic84 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_even_odd <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_local10 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_local2 <0.001 <0.001 1 <0.001 0.0354 0.0277 0.0051 0.0059 0.0158 <0.001 0.0043 <0.001
shuffle_local3 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 0.002 0.0296 0.0111
shuffle_local5 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 0.0018 <0.001 0.0049 0.0026
Table 10: Welch’s t-test comparing each perturbation with shuffle_control across 12 checkpoints for Dutch, with Bonferroni adjustment.
Perturbation | Step 100 200 300 400 500 600 700 800 900 1000 1100 1200
perturb_reverse_full_word <0.001 <0.001 0.0188 0.0016 0.0025 0.0751 0.0028 0.684 0.938 1 0.8737 1
shuffle_deterministic21 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_deterministic57 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_deterministic84 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_even_odd <0.001 <0.001 <0.001 0.0078 <0.001 <0.001 0.0022 0.4015 0.0068 0.0013 0.0235 0.0812
shuffle_local10 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_local2 0.0458 <0.001 <0.001 0.0055 0.0251 1 1 1 1 1 1 1
shuffle_local3 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 0.0269 0.0878
shuffle_local5 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 0.0167 0.1072
Table 11: Welch’s t-test comparing each perturbation with shuffle_control across 12 checkpoints for English, with Bonferroni adjustment.
Perturbation | Step 100 200 300 400 500 600 700 800 900 1000 1100 1200
perturb_reverse_full_word <0.001 0.0284 1 0.008 1 1 1 1 1 1 1 1
shuffle_deterministic21 <0.001 <0.001 <0.001 <0.001 <0.001 0.0046 0.6694 1 1 1 1 1
shuffle_deterministic57 1 <0.001 <0.001 <0.001 <0.001 0.0066 0.6324 1 1 1 1 1
shuffle_deterministic84 <0.001 0.0203 <0.001 <0.001 <0.001 <0.001 0.0402 1 1 1 1 1
shuffle_even_odd <0.001 <0.001 1 <0.001 1 1 1 1 1 1 1 1
shuffle_local10 <0.001 <0.001 0.3162 <0.001 1 1 1 1 1 1 1 1
shuffle_local2 <0.001 <0.001 1 1 1 1 1 1 1 1 1 1
shuffle_local3 <0.001 <0.001 1 0.5046 1 1 1 1 1 1 1 1
shuffle_local5 <0.001 <0.001 1 0.0257 1 1 1 1 1 1 1 1
Table 12: Welch’s t-test comparing each perturbation with shuffle_control across 12 checkpoints for French, with Bonferroni adjustment.
Perturbation | Step 100 200 300 400 500 600 700 800 900 1000 1100 1200
perturb_reverse_full_word <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_deterministic21 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_deterministic57 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_deterministic84 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_even_odd <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_local10 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_local2 <0.001 <0.001 <0.001 <0.001 0.034 0.0974 0.4926 0.4579 1 0.2287 1 0.4976
shuffle_local3 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 0.0035 0.0185 0.011 0.0675 0.14 0.2177
shuffle_local5 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 0.001 <0.001 <0.001 <0.001 0.0017
Table 13: Welch’s t-test comparing each perturbation with shuffle_control across 12 checkpoints for German, with Bonferroni adjustment.
Perturbation | Step 100 200 300 400 500 600 700 800 900 1000 1100 1200
perturb_reverse_full_word <0.001 1 1 1 1 1 1 1 1 1 1 1
shuffle_deterministic21 <0.001 <0.001 1 0.5567 1 1 1 1 1 1 1 1
shuffle_deterministic57 0.002 0.0107 0.8726 0.3123 1 1 1 1 1 1 1 1
shuffle_deterministic84 <0.001 <0.001 0.0112 0.0013 0.002 1 1 1 0.8841 1 1 1
shuffle_even_odd <0.001 <0.001 0.1087 0.3001 0.1957 1 1 1 1 1 1 1
shuffle_local_word3 <0.001 0.0846 1 1 1 1 1 1 1 1 1 1
shuffle_local10 0.0098 <0.001 <0.001 0.4004 1 1 1 1 1 1 1 1
shuffle_local2 <0.001 1 1 0.3393 0.0606 0.1196 0.1088 0.1105 0.1625 0.1634 0.2567 0.2097
shuffle_local3 1 1 1 1 1 1 1 1 1 1 1 1
shuffle_local5 1 1 1 1 1 1 1 1 1 1 1 1
Table 14: Welch’s t-test comparing each perturbation with shuffle_control across 12 checkpoints for Italian, with Bonferroni adjustment.
Perturbation | Step 100 200 300 400 500 600 700 800 900 1000 1100 1200
perturb_reverse_full_word <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_deterministic21 0.0018 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_deterministic57 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_deterministic84 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_even_odd 0.0049 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_local10 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_local2 0.0288 <0.001 <0.001 <0.001 <0.001 0.0166 0.0577 0.1104 0.0993 0.0446 0.0521 0.0508
shuffle_local3 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_local5 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
Table 15: Welch’s t-test comparing each perturbation with shuffle_control across 12 checkpoints for Polish, with Bonferroni adjustment.
Perturbation | Step 100 200 300 400 500 600 700 800 900 1000 1100 1200
perturb_reverse_full_word <0.001 <0.001 <0.001 <0.001 <0.001 0.0056 0.0028 0.0659 0.1986 1 1 1
shuffle_deterministic21 1 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 0.0358 0.0039 0.0778
shuffle_deterministic57 1 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 1 0.4359 1
shuffle_deterministic84 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 0.006 0.0012 0.0137
shuffle_even_odd 1 <0.001 <0.001 0.0569 <0.001 0.3289 1 1 1 1 1 1
shuffle_local10 1 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 0.1379 1 1 1 1
shuffle_local2 1 1 <0.001 1 0.1401 1 1 1 1 1 1 1
shuffle_local3 0.8382 <0.001 <0.001 0.0189 <0.001 1 1 1 1 1 1 1
shuffle_local5 1 <0.001 <0.001 0.6868 <0.001 1 1 1 1 1 1 1
Table 16: Welch’s t-test comparing each perturbation with shuffle_control across 12 checkpoints for Portuguese, with Bonferroni adjustment.
Perturbation | Step 100 200 300 400 500 600 700 800 900 1000 1100 1200
perturb_reverse_full_word <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 1 <0.001 <0.001 0.0693 0.0291 0.0033
shuffle_deterministic21 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 0.6784 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_deterministic57 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 1 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_deterministic84 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 0.1518 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_even_odd <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 1 <0.001 <0.001 0.0059 0.011 0.1607
shuffle_local10 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 1 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_local2 <0.001 <0.001 <0.001 <0.001 <0.001 0.0311 1 0.1407 0.123 0.0092 0.0079 0.0246
shuffle_local3 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 1 <0.001 <0.001 <0.001 <0.001 0.0292
shuffle_local5 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 1 <0.001 <0.001 0.0014 <0.001 0.2879
Table 17: Welch’s t-test comparing each perturbation with shuffle_control across 12 checkpoints for Romanian, with Bonferroni adjustment.
Perturbation | Step 100 200 300 400 500 600 700 800 900 1000 1100 1200
perturb_reverse_full_word <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_deterministic21 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_deterministic57 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_deterministic84 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_even_odd <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_local10 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_local2 <0.001 <0.001 <0.001 1 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_local3 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_local5 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
Table 18: Welch’s t-test comparing each perturbation with shuffle_control across 12 checkpoints for Russian, with Bonferroni adjustment.
Perturbation | Step 100 200 300 400 500 600 700 800 900 1000 1100 1200
perturb_reverse_full_word <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_deterministic21 0.0551 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_deterministic57 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_deterministic84 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_even_odd <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_local10 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_local2 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 0.0036 <0.001 0.002 0.0107 0.0403
shuffle_local3 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_local5 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
shuffle_nondeterministic <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001 <0.001
Table 19: Welch’s t-test comparing each perturbation with shuffle_control across 12 checkpoints for Turkish, with Bonferroni adjustment.